In all of the following answers, enter formulas us
Answer #1

Part A) w = mg

Part B) w_{\perp }= mg\cos \theta

Part C) ma_{\perp }= N-mg\cos \theta = 0\Rightarrow a_{\perp } = 0

Part D) \sum F_{\perp }= N-mg\cos \theta = 0

Part E) N=mg\cos \theta

Part F) w_{//}=mg \sin \theta

Part G) f = \mu N = \mu mg\cos \theta

Part H) \sum F_{//} =mg\sin \theta -f

\sum F_{//} =mg\sin \theta -\mu mg\cos \theta

\sum F_{//} =m(g\sin \theta -\mu g\cos \theta)